OpenStudy (anonymous):

Fans & Medals!!! What is the solution to the system of equations represented by the two equations? y=3/2x y=-1/2x+4

OpenStudy (mimi_x3):

Let multiply the second equation by 3 \(3(y=-\frac{1}{2}x+4) \to 3y=-\frac{3}{2}x+12\) So now we just gotta add the 2 equations together

OpenStudy (tkhunny):

OR, they are both "= y", (3/2)x = (-1/2)x + 4 -- And solve for x There are several ways to go about it.

OpenStudy (mimi_x3):

ummm go with tkhunnys method its easier

OpenStudy (anonymous):

Ok thnx